Reports 1-1 of 1 Clear search Modify search
VIS (General)
satoru.ikeda - 14:58 Friday 21 January 2022 (19525) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

Routine check for DGS maintenance day (1/21)

Stepper motor
  Check if the power supply can be controlled by BIO.
  Check if the script can be started.
  Confirmed that the limit switch installed is in Green.
 [Check result]
  No problem.

 The following settings were made.
  Change the value of software limit of ETMX and ITMX GAS (E2213706-v3)

 The HW limit switches on the ETMX IP were all red because the cable had been unplugged during the leak test, but later checks showed that they were all green, so no problem.

Release Sensor
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMX, ETMY, ITMX, ITMY : No problem

  Ping Test is performed.
 [Check result]
  Known only

   => Confirmation only, no recovery work was done.


  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.

Temperature controller
 Check if the temperature acquisition is working properly.

 [Check result]
  No problem

Precision air processer
 Is the temperature acquisition function working properly?

 [Check result]
  No problem


Images attached to this comment
Search Help